OpenStudy (anonymous):

Solve the system by subsitution x-3y=4 2x-6y=8

OpenStudy (shamil98):

x - 3y = 4 2x - 6y = 8 x = 3y + 4 substitute in the second equation and solve for y, and then plug the value of y back in and evaluate for x.

OpenStudy (anonymous):

^Do that.

OpenStudy (anonymous):

I'll help you further if you don't understand.

OpenStudy (anonymous):

I need a little more help.

OpenStudy (anonymous):

Okay, I'll walk you through it. x = 3y + 4 So in this equation: x - 3y = 4 You replace x with 3y + 4.

OpenStudy (anonymous):

Do you understand now?

OpenStudy (shamil98):

He means in this 2x-6y=8 one you replace the x with 3y + 4

OpenStudy (anonymous):

Okay... I don't think this is right but I got x=4, y=0

OpenStudy (anonymous):

I'll show you.

OpenStudy (anonymous):

x - 3y = 4 (3y + 4) - 3y = 4 3y + 4 - 3y = 4 4 = 4

OpenStudy (anonymous):

2x - 6y = 8 2(3y + 4) - 6y = 8 6y + 4 - 6y = 8 4 = 8 ?? This is false, so it can't be (3y + 4) unless there are no real solutions.

OpenStudy (anonymous):

Oh, okay. I see my mistake. Thanks once again!

OpenStudy (anonymous):

No problem.